OpenStudy (anonymous):

How do you find the appropriate table for the differential equation dy/dx= x/x-2?

OpenStudy (michele_laino):

I can solve your equation

OpenStudy (anonymous):

okay. Is that what I need to do first?

OpenStudy (michele_laino):

yes! I think so! I f we know the formula for y(x), then we can compute all its values, each for each value of x

OpenStudy (anonymous):

I'm not sure how to do this type of problem. I'm use to them having at least one y in them.

OpenStudy (michele_laino):

do you know how to solve your differential equation?

OpenStudy (anonymous):

no, not really how to begin

OpenStudy (michele_laino):

it is very simple, we have to solve this integral: \[y\left( x \right) = \int {\frac{x}{{x - 2}}} \;dx\]

OpenStudy (anonymous):

2ln(x-2)+x+C ?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

so do i just fill in values for x?

OpenStudy (michele_laino):

a better expression is: \[y\left( x \right) = x + \ln \left[ {{{\left( {x - 2} \right)}^2}} \right] + C\] Now we have to keep in mind that the solution y(x) exists for all x such that: \[x \ne 2\]

OpenStudy (anonymous):

its not really working for what i need. I need the values of 0 and .5

OpenStudy (michele_laino):

then we can try to insert into that formula several values of x and consider the corresponding values of y

OpenStudy (michele_laino):

ok! Then we can replace x with 0 and x with 1/2, what values of y do you get?

OpenStudy (michele_laino):

we can neglect the arbitrary constant C

OpenStudy (michele_laino):

what is y(0)=...? what is y(0.5)=...?

OpenStudy (anonymous):

0 was undefined and .5 was 1.3...

OpenStudy (anonymous):

those answers weren't right for what they gave

OpenStudy (michele_laino):

sorry for x=0 we have: y(0) = 0+ ln 4= 2*ln2

OpenStudy (michele_laino):

maybe we have to compute x, when y=0 and y= 0.5?

OpenStudy (anonymous):

zero has to either be undefined or 0 and . 5 has to be -1 or -1/3

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

I'm trying to choose an appropriate value of the arbitrary constant

OpenStudy (michele_laino):

if we set C=-2ln 2, then we get: y(0) = 0

OpenStudy (anonymous):

would the one equal -1/3 at that point?

OpenStudy (michele_laino):

please wait a moment!

OpenStudy (anonymous):

okay thanks for all your help btw!

OpenStudy (michele_laino):

after we set C=-2 ln2= -ln4, we get: \[y\left( x \right) = x + \ln \left[ {{{\left( {x - 2} \right)}^2}} \right] - \ln 4\]

OpenStudy (michele_laino):

and y(0.5) is: \[y(1/2) = \frac{1}{2} + \ln \left[ {{{\left( {\frac{1}{2} - 2} \right)}^2}} \right] - \ln 4 = - 0.076\]

OpenStudy (anonymous):

okay thank you!

OpenStudy (michele_laino):

thank you!